# Peer-to-Peer Basis Risk Management for Renewable Production Parametric Insurance

### v1

@@ 3.1.3 Setting a financial value for lost production
The German power plant database specifies which scheme applies to each facility. Day-ahead market prices were sourced from the ENTSO-E Transparency Platform ENTSO-E (2024).
The feed-in tariff data are more challenging because only the minimum, maximum, and average solar subsidies are available. We use the average, noting that the variance in feed-in tariffs is lower than that observed for day-ahead market prices Bundesnetzagentur (2024).
@@ 5 Conclusion
Beyond the energy sector, the flexibility and modularity of our framework make it relevant for other industries exposed to weather-related uncertainties. In agriculture, for instance, the issue of basis risk has been long acknowledged as a key barrier to the adoption of index-based insurance. One promising avenue involves the development of collective index insurance schemes, where a group of farmers jointly purchases a parametric insurance contract and shares basis risk through informal or contractual within-group transfers. These mechanisms allow the group to pay a single collective premium, receive a single indemnity, and internally reallocate payouts based on individual discrepancies between actual losses and index-based triggers. Such pooling strategies, as studied by Trærup (2012) and Santos et al. (2021), offer a compelling parallel to our peer-to-peer compensation mechanism.
@@ Acknowledgements
Declaration of generative AI in scientific writing
During the writing process of this paper, we used ChatGPT to improve the readibility and language of the manuscript. We carefully reviewed the subsequent content, and take full responsibility and accountability for the content of the publication.
Appendix
@@ Appendix B Approximation of the conditional expectation of the loss
We begin by recalling that if $\mathbf{a}_{i}\neq\mathbf{a}$, then the individual linear score $\mathbf{a}_{i}^{\prime}\mathbf{Y}$ can be rewritten in terms of $Z=\mathbf{a}^{\prime}\mathbf{Y}$. We introduce the following property:
Property B.1
Let $g$ be a real-valued measurable function and $f_{i}(x,y,z)$ be the joint p.d.f. of $(X_{i},\,\mathbf{\alpha}_{i}^{\prime}\mathbf{Y},\,Z)$. Then {adjustwidth}-1cm0cm
$\frac{\partial}{\partial z}\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\bm{|}Z=z\right]$ $=$ $\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)\bm{|}Z=z\right]$
@@
$\frac{\partial^{2}}{\partial z^{2}}\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\bm{|}Z=z\right]$ $=$ $\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial^{2}}{\partial z^{2}}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)\bm{|}Z=z\right]$
$\hskip 20.00003pt+\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\left(\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)\right)^{2}\bm{|}Z=z\right]$
@@
Proof. The first derivative of the conditional variance may be written in the following way:
@@
$=$ $\mathbb{E}\left[X_{i}g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)\bm{|}Z=z\right]$
$\hskip 20.00003pt-\mathbb{E}[\left.X_{i}|Z=z\right.]\mathbb{E}\left[g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)\bm{|}Z=z\right]$
@@
$=$ $\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)\bm{|}Z=z\right]$
@@
$\frac{\partial^{2}}{\partial z^{2}}\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\,\middle|\,Z=z\right]$ $=$ $\mathbb{E}\left[X_{i}g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial^{2}}{\partial z^{2}}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{\prime}\bm{Y}\,|\,z)\,\middle|\,Z=z\right]$
$\hskip 20.00003pt+\mathbb{E}\left[X_{i}g(\bm{\alpha}_{i}^{\prime}\bm{Y})\left(\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{\prime}\bm{Y}\,|\,z)\right)^{2}\,\middle|\,Z=z\right]$
$\hskip 20.00003pt-\mathbb{E}\left[g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{\prime}\bm{Y}\,|\,z)\,\middle|\,Z=z\right]\cdot\frac{\partial}{\partial z}\mathbb{E}\left[X_{i}\,|\,Z=z\right]$
$\hskip 20.00003pt-\mathbb{E}\left[X_{i}\,|\,Z=z\right]\cdot\Bigg{(}\mathbb{E}\left[g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial^{2}}{\partial z^{2}}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{\prime}\bm{Y}\,|\,z)\,\middle|\,Z=z\right]$
$\hskip 40.00006pt+\mathbb{E}\left[g(\bm{\alpha}_{i}^{\prime}\bm{Y})\left(\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{\prime}\bm{Y}\,|\,z)\right)^{2}\,\middle|\,Z=z\right]\Bigg{)}$
@@
$=$ $\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial^{2}}{\partial z^{2}}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{\prime}\bm{Y}\,|\,z)\,\middle|\,Z=z\right]$
$\hskip 20.00003pt+\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\left(\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{\prime}\bm{Y}\,|\,z)\right)^{2}\,\middle|\,Z=z\right]$
@@
$\square$
Rewriting the Linear Score
Specifically,
$\mathbf{a}_{i}^{\prime}\mathbf{Y}\;=\;Z\;+\;\tau_{i}\,\mathbf{\alpha}_{i}^{\prime}\mathbf{Y},$
where
$Z\;=\;\mathbf{a}^{\prime}\mathbf{Y},\quad\tau_{i}\;=\;\bigl{\|}\mathbf{a}_{i}-\mathbf{a}\bigr{\|},\quad\mathbf{\alpha}_{i}^{\prime}\;=\;\frac{(\mathbf{a}_{i}-\mathbf{a})^{\prime}}{\|\mathbf{a}_{i}-\mathbf{a}\|}.$
@@
Following Lemma 2 in Gourieroux et al. (2000), we want to derive an approximation for
$\mathbb{E}[X_{i}\;\mid\;\mathbf{a}_{i}^{\prime}\mathbf{Y}=z].$
Let $f_{i}(x,y,z)$ be the joint p.d.f. of $(X_{i},\,\mathbf{\alpha}_{i}^{\prime}\mathbf{Y},\,Z)$. Then
$\mathbb{E}[X_{i}\;\mid\;\mathbf{a}_{i}^{\prime}\mathbf{Y}=z]\;=\;\frac{\displaystyle\int\!\!\int x\,f_{i}\bigl{(}x,y,\,z-\tau_{i}\,y\bigr{)}\,\mathrm{d}x\,\mathrm{d}y}{\displaystyle\int\!\!\int f_{i}\bigl{(}x,y,\,z-\tau_{i}\,y\bigr{)}\,\mathrm{d}x\,\mathrm{d}y}.$
@@
$=$ $\mathbb{E}\Bigl{[}X_{i}\,(\mathbf{\alpha}_{i}^{\prime}\mathbf{Y})\,\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{\prime}\mathbf{Y},z)\,\bigm{|}\,Z=z\Bigr{]}$
@@
$=$ $\mathbb{E}\Bigl{[}X_{i}\,(\mathbf{\alpha}_{i}^{\prime}\mathbf{Y})^{2}\,\frac{\partial^{2}}{\partial z^{2}}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{\prime}\mathbf{Y},z)\,\bigm{|}\,Z=z\Bigr{]}$
$+\;\mathbb{E}\Bigl{[}X_{i}\,(\mathbf{\alpha}_{i}^{\prime}\mathbf{Y})^{2}\,\Bigl{(}\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{\prime}\mathbf{Y},z)\Bigr{)}^{2}\,\bigm{|}\,Z=z\Bigr{]}$
@@
$=$ $\mathbb{E}\Bigl{[}(\mathbf{\alpha}_{i}^{\prime}\mathbf{Y})\,\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{\prime}\mathbf{Y},z)\,\bigm{|}\,Z=z\Bigr{]}$
@@
$=$ $\mathbb{E}\Bigl{[}(\mathbf{\alpha}_{i}^{\prime}\mathbf{Y})^{2}\,\frac{\partial^{2}}{\partial z^{2}}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{\prime}\mathbf{Y},z)\,\bigm{|}\,Z=z\Bigr{]}$
$+\;\mathbb{E}\Bigl{[}(\mathbf{\alpha}_{i}^{\prime}\mathbf{Y})^{2}\,\Bigl{(}\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{\prime}\mathbf{Y},z)\Bigr{)}^{2}\,\bigm{|}\,Z=z\Bigr{]}$
@@
$=$ $\mathbb{E}\left[X_{i}\,\mathbf{\alpha}_{i}^{\prime}\bm{Y}\,\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{\prime}\bm{Y},z)\,\big{|}\,Z=z\right]$
$\hskip 20.00003pt-\mathbb{E}\left[\mathbf{\alpha}_{i}^{\prime}\bm{Y}\,\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{\prime}\bm{Y},z)\,\big{|}\,Z=z\right]\mathbb{E}[X_{i}|Z=z]$
@@
$\frac{\partial}{\partial z}\mathrm{Cov}\left[X_{i},\bm{\alpha}_{i}^{\prime}\bm{Y|}Z=z\right]$ $=$ $\mathrm{Cov}\left[X_{i},\bm{\alpha}_{i}^{\prime}\bm{Y}\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)\bm{|}Z=z\right]$
@@
$C_{i}\left(z\right)$ $=$ $\mathbb{E}[\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y},z)|Z=z]$
$=$ $\frac{\partial}{\partial z}\log f\left(z\right)\mathbb{E}[\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}|Z=z]+\mathbb{E}[\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}\frac{\partial}{\partial z}\log f_{i}(X_{i},\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)|Z=z]$
$=$ $\frac{\partial}{\partial z}\log f\left(z\right)\mathbb{E}[\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}|Z=z]+\frac{\partial}{\partial z}\mathbb{E}[\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}|Z=z].$
A similar approach applies to $F_{i}(z,\mathbf{a})$, which ultimately involves the second derivatives $\tfrac{\partial^{2}}{\partial z^{2}}\log f_{i}(\dots)$ and mixed covariance terms. Specifically,
$B_{i}(z)-D_{i}(z)\,\mathbb{E}[X_{i}\mid Z=z]$ $=$ $\mathrm{Cov}\Bigl{[}X_{i},\;(\mathbf{\alpha}_{i}^{\prime}\mathbf{Y})^{2}\,\tfrac{\partial^{2}}{\partial z^{2}}\log f_{i}(\dots)\;\bigm{|}\;Z=z\Bigr{]}$
$\hskip 20.00003pt+\mathrm{Cov}\Bigl{[}X_{i},\;(\mathbf{\alpha}_{i}^{\prime}\mathbf{Y})^{2}\,\bigl{(}\tfrac{\partial}{\partial z}\log f_{i}(\dots)\bigr{)}^{2}\;\bigm{|}\;Z=z\Bigr{]}$
@@
$F_{i}\left(z,\bm{a}\right)$ $=$ $\left[\frac{\partial^{2}}{\partial z^{2}}\log f\left(z\right)+\left(\frac{\partial}{\partial z}\log f\left(z\right)\right)^{2}\right]\mathrm{Cov}\left[X_{i},(\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y})^{2}\bm{|}Z=z\right]$
$\hskip 20.00003pt+2\frac{\partial}{\partial z}\log f\left(z\right)\left[\frac{\partial}{\partial z}\mathrm{Cov}\left[X_{i},(\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y})^{2}\bm{|}Z=z\right]+\mathbb{E}[(\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y})^{2}|\left.Z=z\right.]\frac{\partial}{\partial z}\mathbb{E}[\left.X_{i}|Z=z\right.]\right]$
$\hskip 20.00003pt+\frac{\partial^{2}}{\partial z^{2}}\mathrm{Cov}\left[X_{i},(\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y})^{2}\bm{|}Z=z\right]+2\frac{\partial}{\partial z}\mathbb{E}[(\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y})^{2}|\left.Z=z\right.]\frac{\partial}{\partial z}\mathbb{E}[\left.X_{i}|Z=z\right.]$
$\hskip 20.00003pt+\mathbb{E}[(\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y})^{2}|\left.Z=z\right.]\frac{\partial^{2}}{\partial z^{2}}\mathbb{E}[\left.X_{i}|Z=z\right.]$
$\hskip 20.00003pt-2\left[\frac{\partial}{\partial z}\log f\left(z\right)\mathbb{E}[\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}|Z=z]+\frac{\partial}{\partial z}\mathbb{E}[\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}|Z=z]\right]L_{i}\left(z,\bm{a}\right).$
We thus arrive at a fully expanded approximation for $\mathbb{E}[X_{i}\mid\mathbf{a}_{i}^{\prime}\mathbf{Y}=z]$ up to second order in $\tau_{i}$, completing the derivation.

### v2

@@ 3.1.3 Setting a financial value for lost production
The German power plant database specifies which scheme applies to each facility. Day-ahead market prices were sourced from the ENTSO-E Transparency Platform ENTSO-E [25].
The feed-in tariff data are more challenging because only the minimum, maximum, and average solar subsidies are available. We use the average, noting that the variance in feed-in tariffs is lower than that observed for day-ahead market prices Bundesnetzagentur [26].
@@ 5 Conclusion
Beyond the energy sector, the flexibility and modularity of our framework make it relevant for other industries exposed to weather-related uncertainties. In agriculture, for instance, the issue of basis risk has been long acknowledged as a key barrier to the adoption of index-based insurance. One promising avenue involves the development of collective index insurance schemes, where a group of farmers jointly purchases a parametric insurance contract and shares basis risk through informal or contractual within-group transfers. These mechanisms allow the group to pay a single collective premium, receive a single indemnity, and internally reallocate payouts based on individual discrepancies between actual losses and index-based triggers. Such pooling strategies, as studied by Trærup [27] and Santos et al. [28], offer a compelling parallel to our peer-to-peer compensation mechanism.
@@ Appendix B Approximation of the conditional expectation of the loss
We begin by recalling that if $\bm{a}_{i}\neq\bm{a}$, then the individual linear score $\bm{a}_{i}^{\prime}\mathbf{Y}$ can be rewritten in terms of $Z=\bm{a}^{\prime}\mathbf{Y}$.
Rewriting the Linear Score
Specifically,
$\bm{a}_{i}^{\prime}\mathbf{Y}\;=\;Z\;+\;\tau_{i}\,\bm{\alpha}_{i}^{\prime}\mathbf{Y},$
where
$Z\;=\;\bm{a}^{\prime}\mathbf{Y},\quad\tau_{i}\;=\;\bigl{\|}\bm{a}_{i}-\bm{a}\bigr{\|},\quad\bm{\alpha}_{i}^{\prime}\;=\;\frac{(\bm{a}_{i}-\bm{a})^{\prime}}{\|\bm{a}_{i}-\bm{a}\|}.$
We introduce the following property:
Property 1.
Let us assume that $\bm{a}$ and $\bm{a}_{i}$ are not collinear vectors. Let $g$ be a real-valued measurable function and $f_{i}(x,y,z)$ be the joint p.d.f. of $(X_{i},\,\bm{\alpha}_{i}^{\prime}\mathbf{Y},\,Z)$. Then {adjustwidth}-1cm0cm
$\frac{\partial}{\partial z}\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\bm{|}Z=z\right]$ $=$ $\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)\bm{|}Z=z\right]$
@@
$\frac{\partial^{2}}{\partial z^{2}}\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\bm{|}Z=z\right]$ $=$ $\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial^{2}}{\partial z^{2}}\log f_{i}(X_{i},\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)\bm{|}Z=z\right]$
$\hskip 20.00003pt+\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\left(\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)\right)^{2}\bm{|}Z=z\right]$
@@
Proof.
The first derivative of the conditional variance may be written in the following way:
@@
$=$ $\mathbb{E}\left[X_{i}g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)\bm{|}Z=z\right]$
$\hskip 20.00003pt-\mathbb{E}[\left.X_{i}|Z=z\right.]\mathbb{E}\left[g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)\bm{|}Z=z\right]$
@@
$=$ $\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)\bm{|}Z=z\right]$
@@
$\frac{\partial^{2}}{\partial z^{2}}\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\,\middle|\,Z=z\right]$ $=$ $\mathbb{E}\left[X_{i}g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial^{2}}{\partial z^{2}}\log f_{i}(X_{i},\bm{\alpha}_{i}^{\prime}\bm{Y}\,|\,z)\,\middle|\,Z=z\right]$
$\hskip 20.00003pt+\mathbb{E}\left[X_{i}g(\bm{\alpha}_{i}^{\prime}\bm{Y})\left(\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{\prime}\bm{Y}\,|\,z)\right)^{2}\,\middle|\,Z=z\right]$
$\hskip 20.00003pt-\mathbb{E}\left[g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{\prime}\bm{Y}\,|\,z)\,\middle|\,Z=z\right]\cdot\frac{\partial}{\partial z}\mathbb{E}\left[X_{i}\,|\,Z=z\right]$
$\hskip 20.00003pt-\mathbb{E}\left[X_{i}\,|\,Z=z\right]\cdot\Bigg{(}\mathbb{E}\left[g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial^{2}}{\partial z^{2}}\log f_{i}(X_{i},\bm{\alpha}_{i}^{\prime}\bm{Y}\,|\,z)\,\middle|\,Z=z\right]$
$\hskip 40.00006pt+\mathbb{E}\left[g(\bm{\alpha}_{i}^{\prime}\bm{Y})\left(\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{\prime}\bm{Y}\,|\,z)\right)^{2}\,\middle|\,Z=z\right]\Bigg{)}$
@@
$=$ $\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\frac{\partial^{2}}{\partial z^{2}}\log f_{i}(X_{i},\bm{\alpha}_{i}^{\prime}\bm{Y}\,|\,z)\,\middle|\,Z=z\right]$
$\hskip 20.00003pt+\mathrm{Cov}\left[X_{i},g(\bm{\alpha}_{i}^{\prime}\bm{Y})\left(\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{\prime}\bm{Y}\,|\,z)\right)^{2}\,\middle|\,Z=z\right]$
@@
Following Lemma 2 in [29], we want to derive an approximation for
$\mathbb{E}[X_{i}\;\mid\;\bm{a}_{i}^{\prime}\mathbf{Y}=z].$
Let $f_{i}(x,y,z)$ be the joint p.d.f. of $(X_{i},\,\bm{\alpha}_{i}^{\prime}\mathbf{Y},\,Z)$. Then
$\mathbb{E}[X_{i}\;\mid\;\bm{a}_{i}^{\prime}\mathbf{Y}=z]\;=\;\frac{\displaystyle\int\!\!\int x\,f_{i}\bigl{(}x,y,\,z-\tau_{i}\,y\bigr{)}\,\mathrm{d}x\,\mathrm{d}y}{\displaystyle\int\!\!\int f_{i}\bigl{(}x,y,\,z-\tau_{i}\,y\bigr{)}\,\mathrm{d}x\,\mathrm{d}y}.$
@@
$=$ $\mathbb{E}\Bigl{[}X_{i}\,(\bm{\alpha}_{i}^{\prime}\mathbf{Y})\,\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{\prime}\mathbf{Y},z)\,\bigm{|}\,Z=z\Bigr{]}$
@@
$=$ $\mathbb{E}\Bigl{[}X_{i}\,(\bm{\alpha}_{i}^{\prime}\mathbf{Y})^{2}\,\frac{\partial^{2}}{\partial z^{2}}\log f_{i}(X_{i},\bm{\alpha}_{i}^{\prime}\mathbf{Y},z)\,\bigm{|}\,Z=z\Bigr{]}$
$+\;\mathbb{E}\Bigl{[}X_{i}\,(\bm{\alpha}_{i}^{\prime}\mathbf{Y})^{2}\,\Bigl{(}\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{\prime}\mathbf{Y},z)\Bigr{)}^{2}\,\bigm{|}\,Z=z\Bigr{]}$
@@
$=$ $\mathbb{E}\Bigl{[}(\bm{\alpha}_{i}^{\prime}\mathbf{Y})\,\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{\prime}\mathbf{Y},z)\,\bigm{|}\,Z=z\Bigr{]}$
@@
$=$ $\mathbb{E}\Bigl{[}(\bm{\alpha}_{i}^{\prime}\mathbf{Y})^{2}\,\frac{\partial^{2}}{\partial z^{2}}\log f_{i}(X_{i},\bm{\alpha}_{i}^{\prime}\mathbf{Y},z)\,\bigm{|}\,Z=z\Bigr{]}$
$+\;\mathbb{E}\Bigl{[}(\bm{\alpha}_{i}^{\prime}\mathbf{Y})^{2}\,\Bigl{(}\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{\prime}\mathbf{Y},z)\Bigr{)}^{2}\,\bigm{|}\,Z=z\Bigr{]}$
@@
$=$ $\mathbb{E}\left[X_{i}\,\bm{\alpha}_{i}^{\prime}\bm{Y}\,\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{\prime}\bm{Y},z)\,\big{|}\,Z=z\right]$
$\hskip 20.00003pt-\mathbb{E}\left[\bm{\alpha}_{i}^{\prime}\bm{Y}\,\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{\prime}\bm{Y},z)\,\big{|}\,Z=z\right]\mathbb{E}[X_{i}|Z=z]$
@@
$\frac{\partial}{\partial z}\mathrm{Cov}\left[X_{i},\bm{\alpha}_{i}^{\prime}\bm{Y|}Z=z\right]$ $=$ $\mathrm{Cov}\left[X_{i},\bm{\alpha}_{i}^{\prime}\bm{Y}\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)\bm{|}Z=z\right]$
@@
$C_{i}\left(z\right)$ $=$ $\mathbb{E}[\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y}\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y},z)|Z=z]$
$=$ $\frac{\partial}{\partial z}\log f\left(z\right)\mathbb{E}[\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}|Z=z]+\mathbb{E}[\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y}\frac{\partial}{\partial z}\log f_{i}(X_{i},\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y}|z)|Z=z]$
$=$ $\frac{\partial}{\partial z}\log f\left(z\right)\mathbb{E}[\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y}|Z=z]+\frac{\partial}{\partial z}\mathbb{E}[\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y}|Z=z].$
A similar approach applies to $F_{i}(z,\bm{a})$, which ultimately involves the second derivatives $\tfrac{\partial^{2}}{\partial z^{2}}\log f_{i}(\dots)$ and mixed covariance terms. Specifically,
$B_{i}(z)-D_{i}(z)\,\mathbb{E}[X_{i}\mid Z=z]$ $=$ $\mathrm{Cov}\Bigl{[}X_{i},\;(\bm{\alpha}_{i}^{\prime}\mathbf{Y})^{2}\,\tfrac{\partial^{2}}{\partial z^{2}}\log f_{i}(\dots)\;\bigm{|}\;Z=z\Bigr{]}$
$\hskip 20.00003pt+\mathrm{Cov}\Bigl{[}X_{i},\;(\bm{\alpha}_{i}^{\prime}\mathbf{Y})^{2}\,\bigl{(}\tfrac{\partial}{\partial z}\log f_{i}(\dots)\bigr{)}^{2}\;\bigm{|}\;Z=z\Bigr{]}$
@@
$F_{i}\left(z,\bm{a}\right)$ $=$ $\left[\frac{\partial^{2}}{\partial z^{2}}\log f\left(z\right)+\left(\frac{\partial}{\partial z}\log f\left(z\right)\right)^{2}\right]\mathrm{Cov}\left[X_{i},(\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y})^{2}\bm{|}Z=z\right]$
$\hskip 20.00003pt+2\frac{\partial}{\partial z}\log f\left(z\right)\left[\frac{\partial}{\partial z}\mathrm{Cov}\left[X_{i},(\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y})^{2}\bm{|}Z=z\right]+\mathbb{E}[(\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y})^{2}|\left.Z=z\right.]\frac{\partial}{\partial z}\mathbb{E}[\left.X_{i}|Z=z\right.]\right]$
$\hskip 20.00003pt+\frac{\partial^{2}}{\partial z^{2}}\mathrm{Cov}\left[X_{i},(\mathbf{\alpha}_{i}^{{}^{\prime}}\bm{Y})^{2}\bm{|}Z=z\right]+2\frac{\partial}{\partial z}\mathbb{E}[(\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y})^{2}|\left.Z=z\right.]\frac{\partial}{\partial z}\mathbb{E}[\left.X_{i}|Z=z\right.]$
$\hskip 20.00003pt+\mathbb{E}[(\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y})^{2}|\left.Z=z\right.]\frac{\partial^{2}}{\partial z^{2}}\mathbb{E}[\left.X_{i}|Z=z\right.]$
$\hskip 20.00003pt-2\left[\frac{\partial}{\partial z}\log f\left(z\right)\mathbb{E}[\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y}|Z=z]+\frac{\partial}{\partial z}\mathbb{E}[\bm{\alpha}_{i}^{{}^{\prime}}\bm{Y}|Z=z]\right]L_{i}\left(z,\bm{a}\right).$
We thus arrive at a fully expanded approximation for $\mathbb{E}[X_{i}\mid\bm{a}_{i}^{\prime}\mathbf{Y}=z]$ up to second order in $\tau_{i}$, completing the derivation.